# Orthogonal Howe duality and dynamical (split) symmetric pairs

### v2

@@ 1. Introduction
It turns out that, unlike in the case of $(\mathfrak{gl}_{n},\mathfrak{gl}_{m})$-duality, the two sides of orthogonal Howe duality are not symmetric. Rather, the $\mathrm{O}_{m}$-side should be interpreted as the symmetric pair $\mathrm{GL}_{m}^{\theta}\subset\mathrm{GL}_{m}$, where $\theta$ is the involution on $\mathrm{GL}_{m}$ such that $\mathrm{O}_{m}=\mathrm{GL}_{m}^{\theta}$. This is a known phenomenon, especially manifest in the $q$-version of the orthogonal Howe duality, where replacing $\mathfrak{so}_{2n}$ with the Drinfeld–Jimbo quantum group $\mathrm{U}_{q}(\mathfrak{so}_{2n})$ forces the group $\mathrm{O}_{m}$ to be replaced by an $\iota$quantum group [wenzl2020dualities].
To generalize bispectral duality from $(\mathfrak{gl}_{n},\mathfrak{gl}_{m})$ to $(\mathfrak{so}_{2n},\mathrm{O}_{m})$, we first construct a dynamical Weyl group for a split [^2] The associated Satake diagram has all white nodes and no folding. symmetric pair $\mathfrak{k}\subset\mathfrak{g}$. Working over $\mathbb{C}$, this means $\mathfrak{g}$ is simple and $\mathfrak{k}$ is the fixed points of the Chevalley involution swapping generators $e_{\alpha_{i}}\leftrightarrow-f_{\alpha_{i}}$ and acting by $-1$ on $\mathfrak{h}$.
@@
We do not use $\iota$quantum groups in this work. Although we will occasionally discuss $\iota$quantum groups to motivate certain constructions.
@@
In the interest of clarity we choose to not refer to the constructions in this paper as “coideal", “quantum symmetric pair", or "$\iota$quantum". By analogy with integrable systems, we will instead indicate the corresponding constructions with the adjective “boundary” [CherednikQKZ]. For consistency, we will also refer to the “restricted" or “relative" Weyl group of a symmetric pair [Bump-Liegps-book] as the boundary Weyl group.
@@ 1.1. Boundary Weyl group
[…]
@@ 1.2. Boundary dynamical Weyl group
The Lie algebra $\mathfrak{g}=\mathfrak{sl}_{2}$ contains the subalgebra $\mathfrak{k}=\mathfrak{so}_{2}$, generated by $b=f-e$, and the subalgebra $\mathfrak{b}$, generated by $\{h,e\}$. The two bases $\{f,h,e\}$ and $\{b,h,e\}$ correspond to the usual triangular decomposition of $\mathfrak{g}$ and the Iwasawa decomposition of $\mathfrak{g}$ respectively. These two decompositions give two bases for the Verma module $M_{\lambda}:=\mathrm{U}(\mathfrak{g})\otimes_{\mathrm{U}(\mathfrak{b})}\mathbb{C}_{\lambda}$, namely
@@
The first basis is adapted to studying nontrivial $\mathfrak{sl}_{2}$-submodules of $M_{\lambda}$, which are used by Etingof–Varchenko to define the usual Dynamical Weyl group [EtingofVarchenkoDynamicalWeyl]. The second basis makes it clear that restricting $M_{\lambda}$ to $\mathrm{U}(\mathfrak{k})$ yields a free module of rank one, so for any $\mathfrak{k}$-module $X$ we have an isomorphism $\mathrm{Hom}_{\mathfrak{k}}(M_{\lambda}|_{\mathfrak{k}},X)\cong X$. The interplay of the two bases allows us to define the boundary dynamical simple reflection operator.
The main theorem of the first half of our paper can be summarized as follows.
@@
Let $X$ be an integrable $\mathfrak{k}$-module. For each simple root $\alpha_{i}$, there is a boundary dynamical simple reflection operator $A_{X}^{\mathfrak{k},s_{i}}(\lambda)$, such that
$\ldots A^{\mathfrak{k},s_{i}}_{X}(s_{j}s_{i}\cdot\lambda)A^{\mathfrak{k},s_{j}}_{X}(s_{i}\cdot\lambda)A^{\mathfrak{k},s_{i}}_{X}(\lambda)=\ldots A^{\mathfrak{k},s_{j}}_{X}(s_{i}s_{j}\cdot\lambda)A^{\mathfrak{k},s_{i}}_{X}(s_{j}\cdot\lambda)A^{\mathfrak{k},s_{j}}_{X}(\lambda).$
The number of terms on each side the equation is equal to the order of $(s_{i}s_{j})$ in $W$, the Weyl group of $\mathfrak{g}$.
A key feature of the usual dynamical Weyl group is the tensor compatibility with the dynamical fusion operators introduced in [EtingofVarchenkoExchange], see Proposition 6.6. Our proof of Theorem 1.3 relies on a parallel tensor compatibility between the boundary dynamical Weyl group and the boundary fusion operator introduced in [ReshetikhinStokman]. The precise statement appears in Proposition 7.16. Along the way, we compute the boundary fusion operator in particular examples and derive a boundary version of the ABRR equation [ABRR].
@@ 1.3. Simplified series expressions
In Proposition 3.10, we give a new simplified series for the generating boundary Weyl group operators. This is the $q=1$ version of the $\iota$quantum Weyl group operators proposed in [Zhang-Weinan-thesis].
@@ 1.4. Quantum bispectral duality
$\mathcal{R}^{\mathfrak{gl}_{n}}(z_{1}-z_{2})\quad{{\circlearrowright}}\quad\Lambda^{\bullet}(\mathbb{C}^{n}\otimes\mathbb{C}^{2})\quad{{\circlearrowleft}}\quad\mathcal{B}_{\alpha}^{\mathfrak{sl}_{2}}(-z_{1}+z_{2})$
@@
The (extended) Yangian $\mathrm{X}(\mathfrak{so}_{2n})$ acts [^3] It is well known that there is no evaluation homomorphism for the Yangian outside of type $A$. However, see Section 10.5.1. on the spinor representation $\mathfrak{P}_{n}(z)$, for $z\in\mathbb{C}$. Thus, there is a spinorial $R$-matrix [^4] This is the $R$-matrix operator ˇ$\check{\mathcal{R}}^{\mathfrak{so}_{2n}}(z_{1}-z_{2})$ from Definition 10.11, post-composed with the permutation operator. $\mathcal{R}^{\mathfrak{so}_{2n}}(z_{1}-z_{2})$ acting on $\mathfrak{P}_{n}(z_{1})\otimes\mathfrak{P}_{n}(z_{2})$, which as a vector space is $\mathfrak{P}_{n}^{\otimes 2}\cong\mathfrak{P}_{n\cdot 2}$. On the other hand, $\mathfrak{P}_{n\cdot 2}$ is identified with $\Lambda^{\bullet}(\mathbb{C}^{n}\otimes\mathbb{C}^{2})\cong\Lambda^{\bullet}(\mathbb{C}^{2})^{\otimes n}$. Therefore, $\mathfrak{P}_{n\cdot 2}$ carries an action of $\mathfrak{so}_{2}\subset\mathfrak{sl}_{2}$, the associated boundary dynamical Weyl group operator, and the boundary dynamical Weyl group $B$-operator $\mathcal{B}_{\alpha}^{\mathfrak{so}_{2}}(t)$, see Definition 7.19.
@@
$\mathcal{R}^{\mathfrak{so}_{2n}}(z_{1}-z_{2})\quad{{\circlearrowright}}\quad\mathfrak{P}_{n\cdot 2}\quad{{\circlearrowleft}}\quad\mathcal{B}_{\alpha}^{\mathfrak{so}_{2}}(-z_{1}+z_{2})$
@@
$\mathcal{B}^{\mathfrak{so}_{4}}_{\epsilon_{1}+\epsilon_{2}}(-z_{1}+z_{2}-1)\quad{{\circlearrowright}}\quad\mathfrak{P}_{2\cdot m}\quad{{\circlearrowleft}}\quad\mathcal{R}^{\mathfrak{gl}_{m},t_{2}}(z_{1}-z_{2})$
@@ 1.5. Classical bispectral duality
A common philosophy is to degenerate from quantum to classical by replacing difference equations with differential equations. Under this dictionary, the qKZ difference equation is replaced with the classical KZ equation and the related Knizhnik-Zamolodchikov (or KZ) connection from conformal field theory [KnizhnikZamolodchikov]. For a simple Lie algebra $\mathfrak{g}$ and finite dimensional $\mathfrak{g}$-modules $V_{1},\dots,V_{m}$, one obtains the KZ connection on the trivial bundle $V_{1}\otimes\dots\otimes V_{m}$ over $\mathbb{C}^{m}$.
Analogously, one can replace the dynamical difference equations with the Casimir connection (for instance, see [FelderMarkovTarasovVarchenko]). In this case, for any simple Lie algebra $\mathfrak{g}$ and any finite dimensional module $V$, one obtains the Casimir connection on the trivial bundle $V$ on $\mathfrak{h}$.
@@
$\nabla_{\mathfrak{gl}_{n}}^{\mathrm{KZ}}\quad{{\circlearrowright}}\quad\mathfrak{P}_{n\cdot m}\quad{{\circlearrowleft}}\quad\nabla_{\mathfrak{gl}_{m}}^{\mathrm{Cas}}.$
@@
$\nabla_{\mathfrak{so}_{2n}}^{\mathrm{KZ}}\quad{{\circlearrowright}}\quad\mathfrak{P}_{n\cdot m}\quad{{\circlearrowleft}}\quad\nabla_{\mathfrak{gl}_{m}^{\theta}\subset\mathfrak{gl}_{m}}^{\mathrm{bCas}}$
@@
$\nabla_{\mathfrak{so}_{2n}}^{\mathrm{Cas}}\quad{{\circlearrowright}}\quad\mathfrak{P}_{n\cdot m}\quad{{\circlearrowleft}}\quad\nabla_{\mathfrak{gl}_{m}^{\theta}\subset\mathfrak{gl}_{m}}^{\mathrm{bKZ}}$
@@ 1.6. Structure of the paper
In Section 2 we introduce the necessary notations, in particular we introduce split symmetric pairs $\mathfrak{k}\subset\mathfrak{g}$. In Section 3, we recall the braid group action on finite-dimensional $\mathfrak{g}$-modules, then introduce the boundary analog of the braid group action on integrable representations of $\mathfrak{k}$. In Section 4, we recall the dynamical fusion construction of [EtingofVarchenkoExchange], then in Section 5, we introduce and study a boundary analog of dynamical fusion. In Section 6, we recall the construction of the dynamical Weyl group [EtingofVarchenkoDynamicalWeyl], then in Section 7 we study the boundary dynamical Weyl group. In Section 8, we recall some basic facts on connections and introduce a boundary analog of the Casimir connection. In Section 9, we recall the bispectral $(\mathfrak{gl}_{n},\mathfrak{gl}_{m})$-duality. Finally, in Section 10, we study the orthogonal bispectral duality.

### v3

@@ 1. Introduction
[…]
To generalize bispectral duality from $(\mathfrak{gl}_{n},\mathfrak{gl}_{m})$ to $(\mathfrak{so}_{2n},\mathrm{O}_{m})$, we first construct a dynamical Weyl group for a split [^2] The associated Satake diagram has all white nodes and no folding. symmetric pair $\mathfrak{k}\subset\mathfrak{g}$. Working over $\mathbb{C}$, this means $\mathfrak{g}$ is simple and $\mathfrak{k}$ is the fixed points of the Chevalley involution swapping generators $e_{\alpha_{i}}\leftrightarrow-f_{\alpha_{i}}$ and acting by 1-1 on $\mathfrak{h}$.
@@
[…]
@@
[…]
@@ 1.1. Boundary Weyl group
To understand the boundary Weyl group action on representations of $K$ algebraically, the building block is the symmetric pair $\mathfrak{so}_{2}\subset\mathfrak{sl}_{2}$. Here $\mathfrak{so}_{2}$ is the fixed points of the involution of $\mathfrak{sl}_{2}$ which swaps $e$ with f-f and sends $h$ to h-h. Note that $\mathfrak{so}_{2}$ is spanned by the element b:=feb:=f-e. Writing $b|_{\mathbb{C}^{2}}$ to denote $b$ acting in $\mathbb{C}^{2}$, one finds that
@@ 1.2. Boundary dynamical Weyl group
The main purpose of the first part of the paper is the construction of a boundary analog of the usual Dynamical Weyl group defined by Tarasov-Varchenko [EtingofVarchenkoDynamicalWeyl]. The construction is based on the following observation.
The Lie algebra $\mathfrak{g}=\mathfrak{sl}_{2}$ contains the subalgebra $\mathfrak{k}=\mathfrak{so}_{2}$, generated by b=feb=f-e, and the subalgebra $\mathfrak{b}$, generated by $\{h,e\}$. The two bases $\{f,h,e\}$ and $\{b,h,e\}$ correspond to the usual triangular decomposition of $\mathfrak{g}$ and the Iwasawa decomposition of $\mathfrak{g}$ respectively. These two decompositions give two bases for the Verma module $M_{\lambda}:=\mathrm{U}(\mathfrak{g})\otimes_{\mathrm{U}(\mathfrak{b})}\mathbb{C}_{\lambda}$, namely
@@
The first basis is adapted to studying nontrivial $\mathfrak{sl}_{2}$-submodules of $M_{\lambda}$, which are used by Etingof–Varchenko to define the usual Dynamical Weyl group [EtingofVarchenkoDynamicalWeyl]. The second basis makes it clear that restricting $M_{\lambda}$ to $\mathrm{U}(\mathfrak{k})$ yields a free module of rank one, so for any $\mathfrak{k}$-module $X$ we have an isomorphism $\mathrm{Hom}_{\mathfrak{k}}(M_{\lambda}|_{\mathfrak{k}},X)\cong X$.
The interplay of the two bases allows us to define for every simple root $\alpha_{i}$ the boundary dynamical simple reflection operator $A^{\mathfrak{k},s_{i}}_{X}(\lambda)$ acting on any integrable $\mathfrak{k}$-representation $X$. These operators are boundary analogs of the standard dynamical operators $A^{\mathfrak{g},s_{i}}_{V}(\lambda)$ from loc. cit. (denoted by $A_{s_{i},V}(\lambda)$ there) acting on any integrable $\mathfrak{g}$-module $V$. In particular, we show the following.
@@
The operators $A^{\mathfrak{k},s_{i}}_{X}(\lambda)$ satisfy the dynamical braid relation
(1.1) $\ldots A^{\mathfrak{k},s_{i}}_{X}(s_{j}s_{i}\cdot\lambda)A^{\mathfrak{k},s_{j}}_{X}(s_{i}\cdot\lambda)A^{\mathfrak{k},s_{i}}_{X}(\lambda)=\ldots A^{\mathfrak{k},s_{j}}_{X}(s_{i}s_{j}\cdot\lambda)A^{\mathfrak{k},s_{i}}_{X}(s_{j}\cdot\lambda)A^{\mathfrak{k},s_{j}}_{X}(\lambda),$
where the number of factors on each side of the equation is the order of $(s_{i}s_{j})$ in $W$, the Weyl group of $\mathfrak{g}$.
A key feature of the usual dynamical Weyl group is the tensor compatibility with the dynamical fusion operators introduced in [EtingofVarchenkoExchange], see Proposition 6.6. Our proof of (1.1) relies on a parallel tensor compatibility between the boundary dynamical Weyl group and the boundary fusion operator introduced in [ReshetikhinStokman]. The precise statement appears in Proposition 7.16. Along the way, we compute the boundary fusion operator in particular examples and derive a boundary version of the ABRR equation [ABRR].
@@ 1.3. Simplified series expressions
[…]
@@ 1.4. Quantum bispectral duality
$\mathcal{R}^{\mathfrak{gl}_{n}}(z_{1}-z_{2})\quad{{\circlearrowright}}\qquad\Lambda^{\bullet}(\mathbb{C}^{n}\otimes\mathbb{C}^{2})\qquad{{\circlearrowleft}}\qquad\mathcal{B}_{\alpha}^{\mathfrak{sl}_{2}}(-z_{1}+z_{2})$
@@
The (extended) Yangian $\mathrm{X}(\mathfrak{so}_{2n})$ acts [^3] It is well known that there is no evaluation homomorphism for the Yangian outside of type $A$. However, see Section 10.5.1. on the spinor representation $\mathfrak{P}_{n}(z)$, for $z\in\mathbb{C}$. Thus, there is a spinorial $R$-matrix [^4] This is the $R$-matrix operator $\check{\mathcal{R}}^{\mathfrak{so}_{2n}}(z_{1}-z_{2})$ from Definition 10.11, post-composed with the permutation operator. $\mathcal{R}^{\mathfrak{so}_{2n}}(z_{1}-z_{2})$ acting on $\mathfrak{P}_{n}(z_{1})\otimes\mathfrak{P}_{n}(z_{2})$, which as a vector space is $\mathfrak{P}_{n}^{\otimes 2}\cong\mathfrak{P}_{n\cdot 2}$. On the other hand, $\mathfrak{P}_{n\cdot 2}$ is identified with $\Lambda^{\bullet}(\mathbb{C}^{n}\otimes\mathbb{C}^{2})\cong\Lambda^{\bullet}(\mathbb{C}^{2})^{\otimes n}$. Therefore, $\mathfrak{P}_{n\cdot 2}$ carries an action of $\mathfrak{so}_{2}\subset\mathfrak{sl}_{2}$, the associated boundary dynamical Weyl group operator, and the boundary dynamical Weyl group $B$-operator $\mathcal{B}_{\alpha}^{\mathfrak{so}_{2}}(t)$, see Definition 7.19.
@@
$\mathcal{R}^{\mathfrak{so}_{2n}}(z_{1}-z_{2})\quad{{\circlearrowright}}\qquad\mathfrak{P}_{n\cdot 2}\qquad{{\circlearrowleft}}\qquad\mathcal{B}_{\alpha}^{\mathfrak{so}_{2}}(-z_{1}+z_{2})$
@@
$\mathcal{B}^{\mathfrak{so}_{4}}_{\epsilon_{1}+\epsilon_{2}}(-z_{1}+z_{2}-1)\quad{{\circlearrowright}}\qquad\mathfrak{P}_{2\cdot m}\qquad{{\circlearrowleft}}\qquad\mathcal{R}^{\mathfrak{gl}_{m},t_{2}}(z_{1}-z_{2})$
@@ 1.5. Classical bispectral duality
A common philosophy is to degenerate from quantum to classical by replacing difference equations with differential equations. Under this dictionary, the qKZ difference equation is replaced with the classical KZ equation and the related Knizhnik-Zamolodchikov (or KZ) connection from conformal field theory [KnizhnikZamolodchikov]. For a simple Lie algebra $\mathfrak{g}$ and finite dimensional $\mathfrak{g}$-modules $V_{1},\dots,V_{m}$, one obtains the KZ connection on the trivial bundle $V_{1}\otimes\dots\otimes V_{m}$ over $\mathbb{C}^{m}$, with poles on small diagonals.
Analogously, one can replace the dynamical difference equations with the Casimir connection (for instance, see [FelderMarkovTarasovVarchenko]). In this case, for any simple Lie algebra $\mathfrak{g}$ and any finite dimensional module $V$, one obtains the Casimir connection on the trivial bundle $V$ on $\mathfrak{h}$, with poles on root hyperplanes.
@@
(1.2) $\nabla_{\mathfrak{gl}_{n}}^{\mathrm{KZ}}\quad{{\circlearrowright}}\qquad\mathfrak{P}_{n\cdot m}\qquad{{\circlearrowleft}}\qquad\nabla_{\mathfrak{gl}_{m}}^{\mathrm{Cas}}.$
@@
$\nabla_{\mathfrak{so}_{2n}}^{\mathrm{KZ}}\quad{{\circlearrowright}}\qquad\mathfrak{P}_{n\cdot m}\qquad{{\circlearrowleft}}\qquad\nabla_{\mathfrak{gl}_{m}^{\theta}\subset\mathfrak{gl}_{m}}^{\mathrm{bCas}}$
@@
$\nabla_{\mathfrak{so}_{2n}}^{\mathrm{Cas}}\quad{{\circlearrowright}}\qquad\mathfrak{P}_{n\cdot m}\qquad{{\circlearrowleft}}\qquad\nabla_{\mathfrak{gl}_{m}^{\theta}\subset\mathfrak{gl}_{m}}^{\mathrm{bKZ}}$
@@
For any $\mathfrak{g}$, the standard KZ connection is closely related to Gaudin Hamiltonians [GaudinBook] acting on a tensor product of finite dimensional $\mathfrak{g}$-modules $V_{1}\otimes\ldots\otimes V_{m}$. The Gaudin Hamiltonians, and their higher analogs constructed with the Feigin-Frenkel center [FFR], depend on $m$ distinct points $z_{1},\dots,z_{m}$ in $\mathbb{C}$.
The (higher) Gaudin Hamiltonians generate a commutative subalgebra $\mathcal{A}^{\mathfrak{g}}(z_{1},\ldots,z_{m})\subset\mathrm{End}(V_{1}\otimes\ldots\otimes V_{m})$. By [RybnikovShiftOfArgument], this is a particular case of a commutative subalgebra $\mathcal{A}_{\lambda}^{\mathfrak{g}}(z_{1},\ldots,z_{m})$ which also depends on a dynamical parameter $\lambda\in\mathfrak{h}^{*}$. When $m=1$, it gives the other special case: the quantum shift of argument algebra $\mathcal{A}_{\lambda}^{\mathfrak{g}}$, which is generated by purely dynamical Hamiltonians closely related to the Casimir connection for $\mathfrak{g}$.
It turns out that the aforementioned $(\mathfrak{gl}_{n},\mathfrak{gl}_{m})$-duality yields a bispectral duality between the (higher) Gaudin Hamiltonians. Precisely this is the following equality
$\mathcal{A}^{\mathfrak{gl}_{n}}_{(\lambda_{1},\ldots,\lambda_{n})}(z_{1},\ldots,z_{m})=\mathcal{A}^{\mathfrak{gl}_{m}}_{(z_{1},\ldots,z_{m})}(\lambda_{1},\ldots,\lambda_{n})\subset\mathrm{End}(\mathfrak{P}_{nm}),$
generalizing (1.2), see [HuangMukhin] and [MTV] for the symmetric case. Moreover, by loc. cit., this result can be viewed as a generalization of the classical Capelli identity.
Similarly, on the $\mathfrak{so}_{2n}$-side of $(\mathfrak{so}_{2n},\mathrm{O}_{m})$-duality one can consider $\mathcal{A}_{\lambda}^{\mathfrak{so}_{2n}}(z_{1},\dots,z_{n})\subset\mathrm{End}(\mathfrak{P}_{nm})$. Note that $\lambda\in\mathfrak{h}^{*}_{\mathfrak{so}_{2n}}$. We expect that on the $O_{m}$-side of $(\mathfrak{so}_{2n},\mathrm{O}_{m})$-duality this algebra of operators can be generated by Hamiltonians constructed in a way which uses the symmetric pair $\mathrm{O}_{m}\subset\mathrm{GL}_{m}$.
A candidate construction for such operators would come from the cyclotomic analog of the higher Gaudin Hamiltonians defined in [VicedoYoung] and [VicedoYoundIrregularSing]. We conjecture that the cyclotomic Hamiltonians for $\mathrm{O}_{m}\subset\mathrm{GL}_{m}$ coincide, under $(\mathfrak{so}_{2n},\mathrm{O}_{m})$-duality, with the purely dynamical Hamiltonians for $\mathfrak{so}_{2n}$. Some evidence is that the construction of loc. cit. applied to a particular element of the center gives operators closely related to the boundary KZ connection which is dual to the Casimir connection on the $\mathfrak{so}_{2n}$-side by Theorem˜1.10.
However, it is less clear how to construct operators on the $\mathrm{O}_{m}\subset\mathrm{GL}_{m}$-side which correspond to the (higher) Gaudin Hamiltonians on the $\mathfrak{so}_{2n}$-side of $(\mathfrak{so}_{2n},\mathrm{O}_{m})$-duality. A naive guess, based on other bispectral duality results, is that this construction should stem from a boundary analog of the quantum shift of argument algebra $\mathcal{A}^{\mathfrak{k}\subset\mathfrak{g}}_{\lambda}$ where $\lambda\in\mathfrak{h}^{*}$ for $\mathfrak{h}\subset\mathfrak{g}$ a Cartan.
Beyond bispectral duality results, it would be interesting to extend some general results about the usual higher Gaudin Hamiltonians to the cyclotomic setting, and the $(\mathfrak{so}_{2n},\mathrm{O}_{m})$-duality could provide a useful insight into such problems. An instance of a general result is that, under certain conditions on $\lambda$ and $z_{1},\ldots,z_{m}$, the algebra $\mathcal{A}^{\mathfrak{g}}_{\lambda}(z_{1},\ldots,z_{m})$ acts with simple spectrum on the tensor product of irreducible finite-dimensional representations, see [FFR, Corollary 5]. We do not know if there is a similar result for the cyclotomic Gaudin subalgebras. If there is such simple spectrum result, a further line of inquiry would be to relate the cyclotomic Gaudin subalgebras to a boundary analog of crystals [Watanabe] as in [HKRW].
@@ 1.6. Structure of the paper
In Section 2 we introduce the necessary notations, in particular we introduce split symmetric pairs $\mathfrak{k}\subset\mathfrak{g}$. In Section˜3, we recall the braid group action on finite-dimensional $\mathfrak{g}$-modules, then introduce the boundary analog of the braid group action on integrable representations of $\mathfrak{k}$. In Section˜4, we recall the dynamical fusion construction of [EtingofVarchenkoExchange], then in Section˜5, we introduce and study a boundary analog of dynamical fusion. In Section˜6, we recall the construction of the dynamical Weyl group [EtingofVarchenkoDynamicalWeyl], then in Section 7 we study the boundary dynamical Weyl group. In Section˜8, we recall some basic facts on connections and introduce a boundary analog of the Casimir connection. In Section˜9, we recall the bispectral $(\mathfrak{gl}_{n},\mathfrak{gl}_{m})$-duality. Finally, in Section˜10, we study the orthogonal bispectral duality.